Juggernaut Satellite Corporation earned $16 million for the fiscal year ending yesterday. The firm also paid out 30 percent of its earnings as dividends yesterday. The firm will continue to pay out 30 percent of its earnings as annual, end-of-year dividends. The remaining 70 percent of earnings is retained by the company for use in projects. The company has 1.5 million shares of common stock outstanding. The current stock price is $89. The historical return on equity (ROE) of 18 percent is expected to continue in the future.

  

What is the required rate of return on the stock? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  

  Rate of return %
Answer #1

Price=D1/(r-g)
=>Price=D1/(r-ROE*(1-payout rate))
=>r=D1/P+ROE*(1-payout rate)
=>r=D0/P*(1+ROE*(1-payout rate))+ROE*(1-payout rate)
=>r=E0/P*payout*(1+ROE*(1-payout rate))+ROE*(1-payout rate)
=>r=16/1.5*1/89*30%*(1+18%*(1-30%))+18%*(1-30%)
=>r=16.6485%